# Generate, but Verify: Reducing Hallucination in Vision-Language Models with Retrospective Resampling

### v2

@@ Appendix B Dataset Details
In the construction of the dataset, the automatically annotated phrases include all noun phrases (multilingual support) and a set of predefined keywords relevant to the task, such as “Yes” or “No.” We maintain a skip list to prevent the model from marking common but uninformative phrases, like “in the image.” Noun phrases with their corresponding prefix prepositions are automatically extracted using Part-of-Speech (POS) tagging tools [21], and the full data curation pipeline is in Figure A.1. The skip list is in Figure A.3. In this work, we only annotate model response phrases rather than the questions. We also show the prompt for constructing negative phrases for general questions in the following pages.
To further enhance dataset quality and balance, we validate that all tokens are tagged properly and then redistribute positive and negative answers. The final image distribution of our data composition is the same as the original LLaVA-v1.5-665k dataset as shown in Figure A.2.
Appendix C Limitations and Societal Impact
Our 1.3M Dataset:
REVERSE-VLMs:
REVERSE is a generalizable framework applicable to models like LLaVA-v1.5, LLaVA-MORE, and Qwen2.5-VL. It significantly reduces hallucination with minimal loss in expressiveness and efficiency (shown in Table 1, Table 2). However, it does not improve performance on discriminative VQA tasks (see Appendix D) as the backtracking methods generally didn’t help further reasoning. Future work can explore integrating REVERSE with existing de-hallucination techniques to improve such tasks. As a well studied problem, REVERSE carries risks of misuse just like other VLMs [8, 33, 34]. We adopt existing safety mechanisms from upstream models (i.e., LLaVA-v1.5 [34], LLaVA-MORE [16], and Qwen2.5-VL [6]), and provide a fully open-source release with training details.
@@ Appendix C Limitations and Societal Impact
Appendix D Evaluation Datasets & Metrics
@@
Appendix E Additional Results
@@
We report results on discriminative hallucination benchmarks including AMBER-D, POPE, and MME-Hall in Appendix D. These benchmarks consist of binary classification tasks such as Yes or No questions, where the impact of retrospective resampling is naturally limited. Since the answer space is minimal, often restricted to a single token, rethinking the output using a clarification hint like “(potential incorrect phrases Yes/No)” described in subsection 3.3 generally does not provide meaningful improvement.
@@
Appendix F Implementation Details
F.1 Training Details
@@
F.2 Full Decoding Algorithm

### v3

@@ Appendix B Dataset Details
The philosophy of our data generation is to be automatic, scalable, and high-quality. The overall pipeline is illustrated in Figure A.1. The dataset construction begins with the automatic annotation of all noun phrases (with multilingual support) and a set of predefined task-relevant keywords such as “Yes” or “No.” To avoid uninformative annotations (e.g., “in the image”), we maintain a skip list (Figure A.3). Noun phrases and their prefix prepositions are extracted automatically using Part-of-Speech (POS) tagging tools [22]. This stage produces positive examples that accurately describe the visual content (objects). To teach the model to produce unconfident tokens (</UN>) after incorrect, hallucinated objects, we augment each positive example with one corresponding negative sample, generated through a combination of rule-based algorithms and LLM prompting with a strong emphasis on diversity and quality, as follows.
Type-aware rules.
We first categorize QA pairs into different types. For simple cases (e.g., yes/no or numerical questions), we apply deterministic transformations such as flipping answers or sampling plausible but incorrect numbers.
LLM-generated negatives.
For more complex cases (e.g., open-ended descriptive answers), we use GPT-4o to generate diverse and semantically plausible negative responses, following the structured prompts described in the following pages.
Quality control.
To ensure high data quality, we filter out trivial negatives (e.g., those differing only by pronoun swaps) and re-distribute examples to balance the positive/negative ordering, preventing positional bias and ensuring that each negative sample is both meaningful and informative.
The final image distribution matches that of the original LLaVA-v1.5-665k dataset, as shown in Figure A.2. Compared to existing datasets such as FAVA [36], the data used for REVERSE is fundamentally different in both scale and diversity: REVERSE contains 1.3M samples (versus 35K in FAVA) and covers a much wider range of hallucination types across attributes, objects, world entities, and scenes, extending beyond the graph-based replacement frameworks adopted in prior work.
@@
Appendix C Evaluation Datasets & Metrics
@@
| Model | Popular | Adversarial | Random | All |
| LLaVA-v1.5-7B | 86.1 (84.8–87.4) | 84.2 (82.8–85.7) | 87.2 (85.8–88.5) | 85.9 (83.2–88.2) |
| REVERSE | 86.3 (85.0–87.7) | 83.9 (82.4–85.3) | 87.5 (86.3–88.8) | 85.9 (82.8–88.5) |
@@
Appendix D Additional Results
@@
We report results on discriminative hallucination benchmarks including AMBER-D, POPE, and MME-Hall in Table A.1 and Table A.2. These benchmarks consist of binary classification tasks such as Yes or No questions, where the impact of retrospective resampling is naturally limited. Since the answer space is minimal, often restricted to a single token, rethinking the output using a clarification hint like “(potential incorrect phrases Yes/No)” described in subsection 3.3 generally does not provide meaningful improvement.
@@
Results on Common VQA Tasks:
To verify that REVERSE preserves performance on standard VQA benchmarks, we further evaluate it on HallusionBench, GQA, and MM-Vet (Appendix D). For captioning tasks, we also aim to ensure that REVERSE maintains caption quality. AMBER-G already reports a “coverage” metric that measures how many relevant objects are mentioned, alongside the hallucination metric CHAIR. For evaluation on MSCOCO, we additionally report the CLAIR [12] metric, which employs an LLM-as-a-judge and correlates more strongly with human judgments than traditional automatic metrics such as CIDEr and SPICE. CLAIR scores range from 0 to 1, with higher values indicating better caption quality. The results are summarized in Appendix D.
On general VQA benchmarks, REVERSE achieves performance comparable to the LLaVA-v1.5-7B baseline, showing that the gains on hallucination-focused benchmarks (e.g., MM-Hal or Haloquest) do not come at the expense of general VLM capabilities. Although HallusionBench is sometimes considered a visual hallucination benchmark, it primarily evaluates robustness to visually misleading inputs rather than language-prior or bias-driven object hallucination that our work explicitly targets. Overall, REVERSE achieves a 30% reduction in object hallucination while maintaining comparable caption quality to the baseline.
| Method | HallusionBench | GQA | MM-Vet |
| LLaVA-v1.5-7B | 46.94 | 62.00 | 31.10 |
| REVERSE() | 45.44 | 62.73 | 28.40 |
| Method | |||
| LLaVA-v1.5-7B | 15.4 | 50.0 | 0.7384 |
| REVERSE() | 10.3 | 37.0 | 0.7264 |
@@
Appendix E Implementation Details
E.1 Training Details
@@
E.2 Full Decoding Algorithm